You are considering the purchase of a ​$1 comma 000 par value bond with a coupon rate of 5.1​% ​(with interest paid​ semiannually) that matures in 12 years. If the bond is priced to yield 9​%, what is the​ bond's current​ price?
Answer #1

Price of a bond is mathematically represented as:

where P is price of a bond with periodic coupon C, n periods to maturity, periodic YTM i and M face value.

For this bond,

M = $1000, n = 12 * 2 = 24 semi-annual periods, i = 9%/2 = 4.5% (semi-annually), C = 5.1% * $1000/2 = $25.5

P = 369.63 + 347.70

P = $717.34